# LPENSL-TH-08/22

### v2

@@ 2.3 Reformulation of the spectrum and eigenstates
$\mathfrak{f}_{\boldsymbol{\varepsilon}}^{(r)}\equiv\mathfrak{f}_{\boldsymbol{\varepsilon}}^{(r)}(\tau_{+},\tau_{-},\varphi_{+},\varphi_{-},\psi_{+},\psi_{-})=\frac{2\kappa_{+}\kappa_{-}}{\sinh\varsigma_{+}\,\sinh\varsigma_{-}}\,\big[\cosh(\tau_{+}-\tau_{-})\\ -\epsilon_{\varphi_{+}}\epsilon_{\varphi_{-}}\cosh(\epsilon_{\varphi_{+}}\varphi_{+}+\epsilon_{\varphi_{-}}\varphi_{-}+\epsilon_{\psi_{+}}\psi_{+}-\epsilon_{\psi_{-}}\psi_{-}+(N-1-2r)\eta)\big],$ (2.33)
@@
In the formulation of this proposition, the bases (2.21) used in (2.18)-(2.19) can be either the SoV bases constructed via the new SoV approach proposed in [80], or the bases constructed via the generalized Sklyanin’s approach, provided that the latter approach is applicable (we recall that its range of validity is more restricted)
@@
$|\,Q\,\rangle=\mathsf{c}^{(R)}_{M,\text{ref}}\ \underline{\widehat{\mathcal{B}}}_{-,M}(\{\lambda_{i}\}_{i=1}^{M}|\alpha-\beta+1)\,|\,\eta,\alpha+\beta+N-2M-1\,\rangle.$ (2.45)
@@ 3.2 Decomposition of gauged boundary states into bulk ones
$D(-\lambda_{M+1}^{(\sigma)}|\gamma+\delta,\alpha-\beta-1)\,B(\lambda_{M}^{(\sigma)}|\gamma-\delta-1,\alpha-\beta+2).\cdots B(\lambda_{1}^{(\sigma)}|\gamma-\delta-M,\alpha-\beta+M+1)\,|\,\eta,\gamma+\delta+M\,\rangle\\ \overset{\text{Direct Action}}{=}\frac{e^{-(\alpha-\beta+M-1)\eta}-e^{-(\gamma+\delta+M-N)\eta}}{e^{\eta/2}}\,\frac{a(-\lambda_{M+1}^{(\sigma)})\,\sinh\eta(\delta+1)}{\sinh\eta(\delta+M+1)}\prod_{a=1}^{M-1}\frac{\sinh(\lambda_{M}^{(\sigma)}+\lambda_{a}^{(\sigma)}+\eta)}{\sinh(\lambda_{M}^{(\sigma)}+\lambda_{a}^{(\sigma)})}\\ \times B(\lambda_{M}^{(\sigma)}|\gamma-\delta-2,\alpha-\beta+1).\cdots B(\lambda_{1}^{(\sigma)}|\gamma-\delta-M-1,\alpha-\beta+M)\,|\,\eta,\gamma+\delta+M+1\,\rangle$, (3.28)
@@ 4 Action of local operators on boundary separate states
where, the sum is performed over all $\sigma_{j}\in\{+,-\}$ for $j\in\alpha_{+}$, we have defined $\mu_{i}^{\sigma}=\sigma_{i}\mu_{i}$ for $i\in\mathsf{B}_{\boldsymbol{\epsilon,\epsilon^{\prime}}}$, with $\sigma_{i}=1$ if $i>M$, and
@@ 5 Matrix elements of $m$-site local operator strips in the finite chain
$\langle\,\{\lambda\}\,|\,\prod_{n=1}^{m}E_{n}^{\epsilon_{n}^{\prime},\epsilon_{n}}(\xi_{n}|(a_{n},b_{n}),(\bar{a}_{n},\bar{b}_{n}))\,|\,\{\lambda\}\,\rangle=\mathsf{c}^{(R)}_{M,\text{ref}}\ \\ \times\frac{\langle\,Q\,|\,\prod_{n=1}^{m}E_{n}^{\epsilon_{n}^{\prime},\epsilon_{n}}(\xi_{n}|(a_{n},b_{n}),(\bar{a}_{n},\bar{b}_{n}))\,\underline{\widehat{\mathcal{B}}}_{-,M}(\{\lambda_{i}\}_{i=1}^{M}|\alpha-\beta+1)\,|\,\eta,\alpha+\beta+N-2M-1\,\rangle}{\langle\,Q\,|\,Q\,\rangle}$, (5.3)
@@
Let the boundary condition (5.1), for a given $M\in\{1,\ldots,N\}$ and with a given choice of $\epsilon_{\varphi_{+}},\epsilon_{\psi_{+}}\in\{+1,-1\}$, be satisfied, and let $\{\lambda\}\equiv\{\lambda_{1},\ldots,\lambda_{M}\}$ be a solution of the homogeneous Bethe equations under the constraint (5.1). Let $\alpha$ and $\beta$ be given in terms of $\varphi_{-},\psi_{-}$ by (2.43)-(2.44), and let $\boldsymbol{\epsilon}\equiv(\epsilon_{1},\ldots,\epsilon_{m}),\boldsymbol{\epsilon^{\prime}}\equiv(\epsilon_{1}^{\prime},\ldots,\epsilon_{m}^{\prime})\in\{1,2\}^{m}$ satisfying (5.4).
@@
$H_{\{\text{{b}}_{j}\}}(\{\lambda\}|\beta)=\prod_{n=1}^{m}\frac{e^{\eta}}{\sinh(\eta b_{n})}\sum\limits_{\sigma_{\text{{b}}_{j}}}\frac{(-1)^{s}\prod\limits_{i=1}^{m}\sigma_{\text{{b}}_{i}}\prod\limits_{i=1}^{m}\prod\limits_{j=1}^{m}\sinh(\lambda_{\text{{b}}_{i}}^{\sigma}+\xi_{j}+\eta/2)}{\prod\limits_{1\leq i<j\leq m}\sinh(\lambda^{\sigma}_{\text{{b}}_{i}}-\lambda_{\text{{b}}_{j}}^{\sigma}-\eta)\sinh(\lambda_{\text{{b}}_{i}}^{\sigma}+\lambda_{\text{{b}}_{j}}^{\sigma}+\eta)}\\ \times\prod\limits_{p=1}^{s}\bigg\{\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{i_{p}}^{(1)}-\eta(1+b_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{\left(0\right)})\bigg\}\\ \times\!\!\prod\limits_{p=s+1}^{m}\bigg\{\sinh(\lambda_{\text{{b}}_{p}\,}^{\sigma}-\xi_{i_{p}}^{(1)}+\eta(1-\bar{b}_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{(1)}+\eta)\bigg\}\\ \times\prod\limits_{k=1}^{m}\frac{\sinh(\xi_{k}-\epsilon_{\varphi_{-}}\varphi_{-})\cosh(\xi_{k}-\epsilon_{\varphi_{-}}\psi_{-})}{\sinh(\lambda_{\text{{b}}_{k}}^{\sigma}-\epsilon_{\varphi_{-}}\varphi_{-}+\eta/2)\cosh(\lambda_{\text{{b}}_{k}}^{\sigma}-\epsilon_{\varphi_{-}}\psi_{-}+\eta/2)}\ \det_{m}\Omega.$ (5.8)
@@ 6 On correlation functions in the half-infinite chain
$H_{m}(\{\lambda_{j}\}_{j=1}^{M};\{\xi_{k}\}_{k=1}^{m})=\frac{\prod\limits_{j=1}^{m}\prod\limits_{k=1}^{m}\sinh(\lambda_{j}+\xi_{k}+\eta/2)}{\!\!\!\!\prod\limits_{1\leq i<j\leq m}\!\!\!\!\sinh(\lambda_{i}-\lambda_{j}-\eta)\,\sinh(\lambda_{i}+\lambda_{j}+\eta)}\\ \times\prod\limits_{p=1}^{s}\bigg\{\sinh(\lambda_{p}-\xi_{i_{p}}^{(1)}-\eta(1+b_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sin(\lambda_{p}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{p}-\xi_{k}^{(1)}-\eta)\bigg\}\\ \times\!\!\prod\limits_{p=s+1}^{m}\bigg\{\sinh(\lambda_{p}-\xi_{i_{p}}^{(1)}+\eta(1-\bar{b}_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sinh(\lambda_{p}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{p}-\xi_{k}^{(1)}+\eta)\bigg\}\\ \times\prod\limits_{k=1}^{m}\frac{\sinh(\xi_{k}-\epsilon_{\varphi_{-}}\varphi_{-})\cosh(\xi_{k}-\epsilon_{\varphi_{-}}\psi_{-})}{\sinh(\lambda_{\text{{b}}_{k}}^{\sigma}-\epsilon_{\varphi_{-}}\varphi_{-}+\eta/2)\cosh(\lambda_{\text{{b}}_{k}}^{\sigma}-\epsilon_{\varphi_{-}}\psi_{-}+\eta/2)},$ (6.9)
@@
$\mathcal{C}_{\boldsymbol{\xi}}=\mathcal{C}\cup\Gamma(\xi_{1}^{(1)},\ldots,\xi_{m}^{(1)}),$ (6.12)
where $\Gamma(\mu_{1},\ldots,\mu_{p})$ denotes a contour surrounding the points of $\mu_{1},\ldots,\mu_{p}$ with index 1, all other poles of the integrand being outside.

### v3

@@ 2.3 Reformulation of the spectrum and eigenstates
$\mathfrak{f}_{\boldsymbol{\varepsilon}}^{(r)}\equiv\mathfrak{f}_{\boldsymbol{\varepsilon}}^{(r)}(\tau_{+},\tau_{-},\varphi_{+},\varphi_{-},\psi_{+},\psi_{-})=\frac{2\kappa_{+}\kappa_{-}}{\sinh\varsigma_{+}\,\sinh\varsigma_{-}}\,\big[\cosh(\tau_{+}-\tau_{-})\\ -\epsilon_{\varphi_{+}}\epsilon_{\varphi_{-}}\cosh(\epsilon_{\varphi_{+}}\varphi_{+}+\epsilon_{\varphi_{-}}\varphi_{-}+\epsilon_{\psi_{+}}\psi_{+}-\epsilon_{\psi_{-}}\psi_{-}+(N-1-2r)\eta)\big].$ (2.33)
@@
In the formulation of this proposition, the bases (2.21) used in (2.18)-(2.19) can be either the SoV bases constructed via the new SoV approach proposed in [80], or the bases constructed via the generalized Sklyanin’s approach, provided that the latter approach is applicable (we recall that its range of validity is more restricted).
@@
$|\,Q\,\rangle=\mathsf{c}^{(R)}_{M,\text{ref}}\,\mathsf{c}^{(R)}_{Q,\textnormal{ABA}}\ \underline{\widehat{\mathcal{B}}}_{-,M}(\{\lambda_{i}\}_{i=1}^{M}|\alpha-\beta+1)\,|\,\eta,\alpha+\beta+N-2M-1\,\rangle$, (2.45)
in which the explicit expression of $\mathsf{c}^{(R)}_{Q,\textnormal{ABA}}$ is given by (3.61) of [1].
@@ 3.2 Decomposition of gauged boundary states into bulk ones
$D(-\lambda_{M+1}^{(\sigma)}|\gamma+\delta,\alpha-\beta-1)\,B(\lambda_{M}^{(\sigma)}|\gamma-\delta-1,\alpha-\beta+2)\cdots B(\lambda_{1}^{(\sigma)}|\gamma-\delta-M,\alpha-\beta+M+1)\,|\,\eta,\gamma+\delta+M\,\rangle\\ \overset{\text{Direct Action}}{=}\frac{e^{-(\alpha-\beta+M-1)\eta}-e^{-(\gamma+\delta+M-N)\eta}}{e^{\eta/2}}\,\frac{a(-\lambda_{M+1}^{(\sigma)})\,\sinh\eta(\delta+1)}{\sinh\eta(\delta+M+1)}\prod_{a=1}^{M-1}\frac{\sinh(\lambda_{M}^{(\sigma)}+\lambda_{a}^{(\sigma)}+\eta)}{\sinh(\lambda_{M}^{(\sigma)}+\lambda_{a}^{(\sigma)})}\\ \times B(\lambda_{M}^{(\sigma)}|\gamma-\delta-2,\alpha-\beta+1)\cdots B(\lambda_{1}^{(\sigma)}|\gamma-\delta-M-1,\alpha-\beta+M)\,|\,\eta,\gamma+\delta+M+1\,\rangle$, (3.28)
@@ 4 Action of local operators on boundary separate states
where the sum is performed over all $\sigma_{j}\in\{+,-\}$ for $j\in\alpha_{+}$, we have defined $\mu_{i}^{\sigma}=\sigma_{i}\mu_{i}$ for $i\in\mathsf{B}_{\boldsymbol{\epsilon,\epsilon^{\prime}}}$, with $\sigma_{i}=1$ if $i>M$, and
@@ 5 Matrix elements of $m$-site local operator strips in the finite chain
$\langle\,\{\lambda\}\,|\,\prod_{n=1}^{m}E_{n}^{\epsilon_{n}^{\prime},\epsilon_{n}}(\xi_{n}|(a_{n},b_{n}),(\bar{a}_{n},\bar{b}_{n}))\,|\,\{\lambda\}\,\rangle=\mathsf{c}^{(R)}_{M,\text{ref}}\,\mathsf{c}^{(R)}_{Q,\textnormal{ABA}}\\ \times\frac{\langle\,Q\,|\,\prod_{n=1}^{m}E_{n}^{\epsilon_{n}^{\prime},\epsilon_{n}}(\xi_{n}|(a_{n},b_{n}),(\bar{a}_{n},\bar{b}_{n}))\,\underline{\widehat{\mathcal{B}}}_{-,M}(\{\lambda_{i}\}_{i=1}^{M}|\alpha-\beta+1)\,|\,\eta,\alpha+\beta+N-2M-1\,\rangle}{\langle\,Q\,|\,Q\,\rangle}$, (5.3)
@@
Let the boundary condition (5.1), for a given $M\in\{1,\ldots,N\}$ and with a given choice of $\epsilon_{\varphi_{+}},\epsilon_{\varphi_{-}}\in\{+1,-1\}$, be satisfied, and let $\{\lambda\}\equiv\{\lambda_{1},\ldots,\lambda_{M}\}$ be a solution of the homogeneous Bethe equations under the constraint (5.1). Let $\alpha$ and $\beta$ be given in terms of $\varphi_{-},\psi_{-}$ by (2.43)-(2.44), and let $\boldsymbol{\epsilon}\equiv(\epsilon_{1},\ldots,\epsilon_{m}),\boldsymbol{\epsilon^{\prime}}\equiv(\epsilon_{1}^{\prime},\ldots,\epsilon_{m}^{\prime})\in\{1,2\}^{m}$ satisfying (5.4).
@@
$H_{\{\text{{b}}_{j}\}}(\{\lambda\}|\beta)=\prod_{n=1}^{m}\frac{e^{\eta}}{\sinh(\eta b_{n})}\sum\limits_{\sigma_{\text{{b}}_{j}}}\frac{(-1)^{s}\prod\limits_{i=1}^{m}\sigma_{\text{{b}}_{i}}\prod\limits_{i=1}^{m}\prod\limits_{j=1}^{m}\sinh(\lambda_{\text{{b}}_{i}}^{\sigma}+\xi_{j}+\eta/2)}{\prod\limits_{1\leq i<j\leq m}\sinh(\lambda^{\sigma}_{\text{{b}}_{i}}-\lambda_{\text{{b}}_{j}}^{\sigma}-\eta)\sinh(\lambda_{\text{{b}}_{i}}^{\sigma}+\lambda_{\text{{b}}_{j}}^{\sigma}+\eta)}\\ \times\prod\limits_{p=1}^{s}\bigg\{\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{i_{p}}^{(1)}-\eta(1+b_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{\left(0\right)})\bigg\}\\ \times\!\!\prod\limits_{p=s+1}^{m}\bigg\{\sinh(\lambda_{\text{{b}}_{p}\,}^{\sigma}-\xi_{i_{p}}^{(1)}+\eta(1-\bar{b}_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{\text{{b}}_{p}}^{\sigma}-\xi_{k}^{(1)}+\eta)\bigg\}\\ \times\prod\limits_{k=1}^{m}\frac{\sinh(\xi_{k}-\epsilon_{\varphi_{-}}\varphi_{-})\cosh(\xi_{k}+\epsilon_{\varphi_{-}}\psi_{-})}{\sinh(\lambda_{\text{{b}}_{k}}^{\sigma}-\epsilon_{\varphi_{-}}\varphi_{-}+\eta/2)\cosh(\lambda_{\text{{b}}_{k}}^{\sigma}+\epsilon_{\varphi_{-}}\psi_{-}+\eta/2)}\ \det_{m}\Omega.$ (5.8)
@@ 6 On correlation functions in the half-infinite chain
$H_{m}(\{\lambda_{j}\}_{j=1}^{M};\{\xi_{k}\}_{k=1}^{m})=\frac{\prod\limits_{j=1}^{m}\prod\limits_{k=1}^{m}\sinh(\lambda_{j}+\xi_{k}+\eta/2)}{\!\!\!\!\prod\limits_{1\leq i<j\leq m}\!\!\!\!\sinh(\lambda_{i}-\lambda_{j}-\eta)\,\sinh(\lambda_{i}+\lambda_{j}+\eta)}\\ \times\prod\limits_{p=1}^{s}\bigg\{\sinh(\lambda_{p}-\xi_{i_{p}}^{(1)}-\eta(1+b_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sin(\lambda_{p}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{p}-\xi_{k}^{(1)}-\eta)\bigg\}\\ \times\!\!\prod\limits_{p=s+1}^{m}\bigg\{\sinh(\lambda_{p}-\xi_{i_{p}}^{(1)}+\eta(1-\bar{b}_{i_{p}}))\prod\limits_{k=1}^{i_{p}-1}\sinh(\lambda_{p}-\xi_{k}^{(1)})\prod\limits_{k=i_{p}+1}^{m}\!\!\sinh(\lambda_{p}-\xi_{k}^{(1)}+\eta)\bigg\}\\ \times\prod\limits_{k=1}^{m}\frac{\sinh(\xi_{k}-\epsilon_{\varphi_{-}}\varphi_{-})\cosh(\xi_{k}+\epsilon_{\varphi_{-}}\psi_{-})}{\sinh(\lambda_{\text{{b}}_{k}}^{\sigma}-\epsilon_{\varphi_{-}}\varphi_{-}+\eta/2)\cosh(\lambda_{\text{{b}}_{k}}^{\sigma}+\epsilon_{\varphi_{-}}\psi_{-}+\eta/2)},$ (6.9)
@@
$\mathcal{C}_{\boldsymbol{\xi}}=\mathcal{C}\cup\Gamma^{-}(\xi_{1}^{(1)},\ldots,\xi_{m}^{(1)}),$ (6.12)
where $\Gamma^{-}(\mu_{1},\ldots,\mu_{p})$ denotes a contour surrounding the points of $\mu_{1},\ldots,\mu_{p}$ with index 1-1, all other poles of the integrand being outside.